# Measuring Star Formation Histories from Asymptotic Giant Branch Stars: A Demonstration in M31

### v2

@@ Abstract
We demonstrate how near infrared (NIR) imaging of resolved luminous asymptotic giant branch (AGB) stars can be used to measure well-constrained star formation histories (SFHs) across cosmic time. Using UKIRT and -band imaging of M31, we first show excellent agreement over the past Gyr between the PHAT SFH of M31’s outer disk derived from a deep optical color-magnitude diagram (CMD; stars with ), and our spatially-matched SFH based only on modeling AGB stars on a NIR CMD ( stars with ). We find that only hundreds of AGB stars are needed for reliable SFH recovery, owing to their excellent age sensitivity in the NIR. We then measure the spatially resolved SFH of M31’s inner stellar halo ( kpc) using AGB stars. We find: (i) a dominant burst of star formation across M31’s inner stellar halo from Gyr ago and lower level, spatially distributed star formation Gyr ago; (ii) a younger ‘quenching time’ in the vicinity of NGC 205 ( Gyr ago) than near M32 ( Gyr ago); (iii) formed over the past Gyr. We discuss some caveats and the enormous potential of resolved AGB stars in the NIR for measuring SFHs back to ancient epochs ( Gyr ago) in galaxies to large distances ( Mpc) with JWST, Roman, and Euclid.
@@ 1 Introduction
Asymptotic giant branch (AGB) stars represent the final nuclear-fusing evolutionary phase for low and intermediate-mass (), intermediate-age ( Myr to Gyr ago) stars. With peak spectral energy distributions at m, they are among the brightest stars in the near-infrared (NIR; Maraston et al. 2006; Melbourne et al. 2012; Melbourne & Boyer 2013, and are known to exist in abundance in a wide variety of galaxies from low-metallicity dwarf galaxies to high-metallicity massive ellipticals (e.g., Gallart et al. 1996; Dalcanton et al. 2012a; Rejkuba et al. 2022; Anand et al. 2024a, b; Hoyt et al. 2024; Lee et al. 2024). Furthermore, there exists a clear mapping between some of their observable properties (i.e., NIR colors and luminosities) and their ages (e.g., Gallart et al., 2005).
@@
However, despite being ubiquitous, observational challenges and the physical complexity of AGB star physics have resulted in them playing a limited role in star formation history (SFH) measurements from optical resolved star color-magnitude diagrams (CMDs) in favor of better understood phases of evolution (e.g., main sequence turnoff, sub-giant branch, red clump; Gallart et al. 2005). In terms of observations, in the optical it is challenging to cleanly differentiate between AGB and red giant branch (RGB) stars, which are far less sensitive to age, as RGB stars with different ages can occupy similar regions of the CMD. From a physics standpoint, accurately modeling the color and magnitudes of AGB stars on a galaxy’s CMD requires a detailed understanding of complex processes in the interiors and atmospheres of AGB stars such as convection, overshooting, hot bottom burning, mass loss, dredge up, and pulsation (Gallart et al., 2005; Marigo et al., 2017; Pastorelli et al., 2019, 2020). The comparatively simple physics for other age-sensitive populations (MSTO, SGB, RC) has made them the staple of accurate SFH modeling despite their less favorable observational properties, such as faintness and increased susceptibility to crowding (e.g., Gallart et al. 2005).
@@
Nevertheless, a handful of pioneering studies recognized the potential of AGB stars for age determinations. Gallart et al. (1996) was the first to suggest AGB stars as a means of measuring a SFH through CMD modeling, noting: “…the potential power of AGBs as tracers of the SFH at intermediate ages ( Gyrs). It can be reasonably be expected that, in the near future when uncertainties in the stellar evolutionary models for AGBs have been overcome, AGBs will provide strong, detailed insight into the SFH for that time interval.” Since then, NIR observations of AGB stars have been used to map spatially resolved stellar ages in nearby galaxies (e.g., Frogel & Blanco, 1983; Reid & Mould, 1984; Wood et al., 1985; Costa & Frogel, 1996; Davidge, 2003, 2005; Cioni et al., 2006, 2008; Melbourne et al., 2010; Jung et al., 2012; Crnojević et al., 2013; Davidge, 2014; McQuinn et al., 2017). However, many of these studies relied on fairly rudimentary AGB star models and, in many cases, undertook simplified statistical approaches that did not capture the full complexity of CMD modeling or only pursued qualitative analysis. For example, Javadi et al. (2011) began an ambitious observational campaign to measure the SFHs of nearby galaxies by comparing the J and K band luminosity functions of their AGB populations to isochrones to estimate their initial stellar masses, and therefore ages. For computational tractability, they adopted simplifying assumptions like a single metallicity and did not use artificial star tests (ASTs). This methodology has been extended to a number of subsequent studies that highlight the scientific power of AGB stars in the NIR (e.g., Rezaeikh et al. 2014; Hamedani Golshan et al. 2017).
@@
A more recent approach has been to compute the ratio of AGB to RGB stars in optical imaging to estimate the quenching epoch (i.e., the age of the youngest stellar population) using theoretical models or empirical relations derived from the SFHs of Local Group (LG) galaxies with deep CMDs (e.g., Rejkuba et al., 2022; Harmsen et al., 2023). Its application to several nearby galaxies both showcases the broad strengths of AGB stars for measuring important star formation timescales in galaxies outside the LG, while also highlighting limitations of AGB stars on optical CMDs.
@@
In this paper, we combine NIR imaging, significant improvements in theoretical AGB stars models, and well-established CMD fitting techniques to measure quantitative, ‘non-parametric’ SFHs using only AGB stars. The increase in high-quality NIR resolved star photometry, in concert with substantial theoretical efforts, particularly by the PARSEC-COLIBRI group, have resulted in a new generation of AGB star models that are anchored to high quality datasets (Marigo et al., 2017; Pastorelli et al., 2019, 2020) and contemporary stellar physics. These models employ far more sophisticated physics than their predecessors, such as the seminal Padova models (Marigo & Girardi, 2007; Marigo et al., 2008), which were the first to include a range of physical processes during the thermal pulsating (TP) AGB phase. These effects include the third dredge-up, hot-bottom burning, low-temperature opacity changes, C-to-M star transitions, mass loss, and pulsation. More recently, the Padova group introduced the COLIBRI code in Marigo et al. (2013), which built upon and eventually succeeded the Padova isochrones for AGB stars. The COLIBRI models were further improved using NIR resolved star observations of AGB stars in nearby dwarf galaxies (Rosenfield et al., 2014, 2016; Marigo et al., 2017). TP-AGB lifetimes were even more tightly constrained using deep NIR imaging imaging of the Magellanic Clouds from the VMC/VISTA survey, resulting in the current COLIBRI TP-AGB star models that impressively reproduce a number of observational constraints (Pastorelli et al., 2019, 2020).
@@
Here, we take the next step in SFH modeling by combining the COLIBRI AGB star models with the well-tested CMD modeling framework MATCH (Dolphin, 2002) in order to measure quantitative SFHs exclusively from NIR observations of AGB stars. MATCH is a widely used, mature CMD fitting code that has been used to measure SFHs for galaxies from resolved star CMDs (e.g., Skillman et al., 2003; Williams et al., 2009; McQuinn et al., 2010; Monelli et al., 2010; Weisz et al., 2011; Cole et al., 2014; Weisz et al., 2014; Lewis et al., 2015; McQuinn et al., 2015; Skillman et al., 2017; Williams et al., 2017; Savino et al., 2023; Weisz et al., 2023a; McQuinn et al., 2024a, b, c). Though the incorporation of the COLIBRI models into MATCH is challenging (e.g., because of the need to correctly accounting for self-extinction in AGB stars), it provides several important benefits including a well-vetted framework for CMD analysis, robust uncertainty determination, and uniformity with MATCH-based SFHs of nearby galaxies based on optical CMDs, which we use to validate the reliability of our AGB star-based SFHs.
@@
Using this newly established framework, we take several steps to illustrate the technique in practice. We first fit mock data to illustrate self-consistency in our modeling. We then measure the SFH of M31’s outer disk using ground-based NIR observations. This region was selected to overlap with the PHAT survey (Dalcanton et al., 2012b; Williams et al., 2014, 2023), which measured the spatially resolved SFH of the same region using deep HST optical CMDs (Williams et al., 2017, hereafter W17), providing a useful comparison point for our AGB star-only based SFHs. Finally, we demonstrate the scientific potential of this new method by measuring the spatially resolved SFH of M31’s inner stellar halo using only ground-based NIR CMDs of AGB stars.
@@
Throughout this paper, we adopt the following M31 geometric parameters used by the PHAT team: center R.A. and Declination = (, ) and (McConnachie et al., 2005; Barmby et al., 2006). We assume the inclination angle of M31 to be (Nieten et al., 2006). Unless otherwise noted, we use projected distances in M31.
@@ 2 AGB Theoretical Isochrones
Young AGB stars are typically the more luminous cool stars on the HR diagram. Though they become less luminous with age, even the 10 Gyr AGB stars are generally more luminous than the TRGB, an important observational property. Moreover, unlike the RGB, the change in AGB age as a function of luminosity (i.e., the age gradient on the HR diagram) is much more pronounced for AGB stars than RGB stars.
@@
The age range we plot is for illustrative purposes only. AGB stars can exist down to Myr in age; below this age, stars do not progress through the AGB star phase. For the very oldest ages ( Gyr ago), age information comes primarily from E-AGB stars, as lower mass stars generally do not go through the TP-AGB phase. As shown in Figure 1, there is decreased separation between between the RGB, E-AGB, and TP-AGB (when present) phases at older ages. In principle, it it possible to measure SFHs using only AGB stars back to Gyr ago, but such ancient time modeling needs to be empirically tested and is beyond the scope of this paper.
Figure 2 shows a NIR CMD, containing stars with a variety of ages at a fixed metallicity of [M/H] dex. A modest amount of simulated noise has been added to the CMD to illustrate the AGB stars of different ages are readily distinguishable in a NIR CMD. Specifically, we generated and applied ASTs with a 50% completeness level at mag. The location of AGB stars in the color combination is highlighted by the grey box, which approximately denotes the location of the AGB stars as defined in the COLIBRI stellar isochrones. These same AGB stars are not easily distinguished from RGB stars at optical wavelengths, as we illustrate in later in the context of M31. Even in the red-optical (e.g., 8000 Å), the SEDs of AGB stars are largely indistinguishable from RGB stars; shifting to even bluer wavelengths makes separation of RGB and AGB stars less tenable. The m wavelength range appears optimal for isolating AGB stars on a CMD.
There are clearly fewer AGB stars than RGB stars. Marigo et al. (2017) predict that AGB stars form per unit stellar mass. Thus, while AGB stars far are more age-sensitive than RGB stars and much brighter than comparably age-sensitive MSTO and SGB stars of the same age, there are fewer of them. We estimate the minimum number of AGB stars for accurate CMD modeling in Appendix A. Finally, we note that ‘extreme’ AGB stars (i.e., the reddest AGB stars), which suffer from particularly complex physics (e.g., rapid and irregular mass-loss, circumstellar extinction), are so rare that they do not meaningfully contribute to the SFH determinations we describe in this paper.
In Figure 3 we show the behavior of AGB stars with a range of metallicities at a fixed age of 1 Gyr. Though there is clearly some dependence on metallicity, the influence of age is much more pronounced. Figure 10, and an associated discussion, in Marigo et al. (2017) provides a deeper demonstration of this point.
@@ 3 Methodology
To fit the observed CMDs, we employ the widely-used CMD fitting package MATCH (version 2.7, Dolphin 2002, 2012, 2013). Given a set of stellar models, a specified initial mass function (IMF), and a unresolved binary fraction, MATCH creates and linearly combines simple stellar population (SSP) models to create a composite CMD. It then convolves the composite CMD with a noise models (i.e., bias, scatter, completeness) from ASTs to create a model CMD with the same observational effects as the data. The model and observed CMD are converted into Hess diagrams (i.e., a binned CMD) and compared using a Poisson likelihood function. MATCH then iterates through various weights of the SSP basis functions until the overall likelihood function is maximized. The synthetic model CMD that best matches the observed CMD is used as the best fit SFH. Random uncertainties on the SFH are computed using a Markov chain Monte Carlo approach, as we describe below.
@@
Starting from the COLIBRI models, we generate a grid of SSP models, each with a total stellar mass of (which yields AGB stars), spanning a range of ages and metallicities. Our SSP models have ages from dex (100 Myr to 7.9 Gyr in linear time), logarithmically spaced by 0.1 dex. In principle, we could generate AGB stars models back to 14 Gyr ago for SFH determination. However, we stop at a younger lookback time to match the PHAT SFH we use to empirically validate our method, as described in § 5.2. We adjust the metallicity range of our grid depending on the considered dataset (see § 5.2 and 5.3).
@@
We generate the SSPs using a Kroupa IMF (Kroupa, 2001) and assume only single stars. The latter is generally a poor assumption of most parts of the CMD. However, a two-AGB binary system is unlikely, owing to their rarity, and their high luminosities, particularly in the NIR, mean that other binary combinations (e.g., AGB-MS star binary) do not significantly affect their colors and luminosities.
@@
| Stellar Density | J | K | Number of | Regions |
@@
| 0 – 6 | 19.4 | 18.4 | 183605 | 1, 2, 9, 10, 12, 13, 14, 15, 16 |
| 6 – 10 | 19.3 | 18.4 | 477373 | 3, 11, 17, 18 |
| 10 – 13 | 19.3 | 18.4 | 881304 | 7 |
| 13 – 16 | 19.3 | 18.4 | 550815 | 4, 8 |
| 16 – 19 | 19.2 | 18.4 | 587536 | 5, 6 |

### v3

# Measuring Star Formation Histories from Asymptotic Giant Branch Stars I: A Demonstration in M31

@@ Abstract
We demonstrate how near-infrared (NIR) imaging of resolved luminous asymptotic giant branch (AGB) stars can be used to measure well-constrained star formation histories (SFHs) across cosmic time. Using UKIRT and -band imaging of M31, we first show excellent agreement over the past Gyr between the PHAT SFH of M31’s outer disk derived from a deep optical color-magnitude diagram (CMD; stars with ), and our spatially-matched SFH based only on modeling AGB stars on a NIR CMD ( stars with ). We find that only AGB stars are needed for reliable SFH recovery, owing to their excellent age sensitivity in the NIR. We then measure the spatially-resolved SFH of M31’s inner stellar halo ( kpc) using AGB stars. We find: (i) a dominant burst of star formation across M31’s stellar halo Gyr ago and lower level, spatially distributed star formation Gyr ago; (ii) formed over the past Gyr. We discuss some caveats and the enormous potential of resolved AGB stars in the NIR for measuring SFHs back to ancient epochs ( Gyr ago) in galaxies to large distances ( Mpc) with JWST, Roman, and Euclid.
@@ 1 Introduction
Asymptotic giant branch (AGB) stars represent the final nuclear-fusing evolutionary phase for low and intermediate-mass (), intermediate-age ( Myr to Gyr ago) stars. With peak spectral energy distributions at m, they are among the brightest stars in the near-infrared (NIR; Maraston et al. 2006; Melbourne et al. 2012; Melbourne & Boyer 2013), and are known to exist in abundance in a wide variety of galaxies from low-metallicity dwarf galaxies to high-metallicity massive ellipticals (e.g., Gallart et al. 1996; Dalcanton et al. 2012a; Rejkuba et al. 2022; Anand et al. 2024a, b; Hoyt et al. 2024; Lee et al. 2024). Furthermore, there exists a clear mapping between some of their observable properties (i.e., NIR colors and luminosities) and their ages (e.g., Gallart et al., 2005).
@@
However, despite being ubiquitous, observational challenges and the physical complexity of AGB star physics have resulted in them playing a limited role in star formation history (SFH) measurements from optical resolved star color-magnitude diagrams (CMDs) in favor of better understood phases of evolution (e.g., main sequence turnoff, sub-giant branch, red clump; Gallart et al. 2005). In terms of observations, in the optical it is challenging to cleanly differentiate between AGB and red giant branch (RGB) stars, which are far less sensitive to age, as RGB stars with different ages can occupy similar regions of the CMD. From a physics standpoint, accurately modeling the color and magnitudes of AGB stars on a galaxy’s CMD requires a detailed understanding of complex processes in the interiors and atmospheres of AGB stars such as convection, overshooting, hot bottom burning, mass loss, dredge up, and pulsation (Gallart et al., 2005; Marigo et al., 2017; Pastorelli et al., 2019, 2020). The comparatively historically better-understood physics for other age-sensitive populations, like the main sequence turn-off (MSTO), subgiant branch (SGB), and red clump (RC), has made them the staple of accurate SFH modeling despite their less favorable observational properties, such as faintness and increased susceptibility to crowding (e.g., Gallart et al. 2005).
@@
Nevertheless, a handful of pioneering studies recognized the potential of AGB stars for age determinations. Gallart et al. (1996) was the first to suggest AGB stars as a means of measuring a SFH through CMD modeling, noting: “…the potential power of AGBs as tracers of the SFH at intermediate ages ( Gyrs). It can reasonably be expected that, in the near future when uncertainties in the stellar evolutionary models for AGBs have been overcome, AGBs will provide strong, detailed insight into the SFH for that time interval.” Since then, NIR observations of AGB stars have been used to map spatially resolved stellar ages in nearby galaxies (e.g., Frogel & Blanco, 1983; Reid & Mould, 1984; Wood et al., 1985; Costa & Frogel, 1996; Davidge, 2003, 2005; Cioni et al., 2006, 2008; Melbourne et al., 2010; Jung et al., 2012; Crnojević et al., 2013; Davidge, 2014; McQuinn et al., 2017). However, many of these studies relied on fairly rudimentary AGB star models and, in many cases, undertook simplified statistical approaches that did not capture the full complexity of CMD modeling or only pursued qualitative analysis. For example, Javadi et al. (2011) began an ambitious observational campaign to measure the SFHs of nearby galaxies by comparing the J and K band luminosity functions of their AGB populations to isochrones to estimate their initial stellar masses, and therefore ages. For computational tractability, they adopted simplifying assumptions like a single metallicity and did not use artificial star tests (ASTs). This methodology has been extended to a number of subsequent studies that highlight the scientific power of AGB stars in the NIR (e.g., Rezaeikh et al. 2014; Hamedani Golshan et al. 2017).
@@
A more recent approach has been to compute the ratio of AGB to RGB stars in optical imaging to estimate the quenching epoch (i.e., the age of the youngest stellar population) using theoretical models or empirical relations derived from the SFHs of Local Group (LG) galaxies with deep CMDs (e.g., Rejkuba et al., 2022; Harmsen et al., 2023; Velguth et al., 2024). Its application to several nearby galaxies both showcases the broad strengths of AGB stars for measuring important star formation timescales in galaxies outside the LG, while also highlighting limitations of AGB stars on optical CMDs.
@@
In this paper, we combine NIR imaging, significant improvements in theoretical AGB stars models, and well-established CMD fitting techniques to measure quantitative, ‘non-parametric’ SFHs using only AGB stars. The increase in high-quality NIR resolved star photometry, in concert with substantial theoretical efforts, particularly by the PARSEC-COLIBRI group, have resulted in a new generation of AGB star models that are anchored to high quality datasets (Marigo et al., 2017; Pastorelli et al., 2019, 2020) and contemporary stellar physics. These models employ far more sophisticated physics than their predecessors, such as the seminal Padova models (Marigo & Girardi, 2007; Marigo et al., 2008), which were the first to include a range of physical processes during the thermal pulsating (TP) AGB phase. These effects include the third dredge-up, hot-bottom burning, low-temperature opacity changes, C-to-M star transitions, mass loss, and pulsation. More recently, the Padova group introduced the COLIBRI code in Marigo et al. (2013), which built upon and eventually succeeded the Padova isochrones for AGB stars. The COLIBRI models were further improved using NIR resolved star observations of AGB stars in nearby dwarf galaxies (Rosenfield et al., 2014, 2016; Marigo et al., 2017). TP-AGB lifetimes were even more tightly constrained using deep NIR imaging of the Magellanic Clouds from the VMC/VISTA survey, resulting in the current COLIBRI TP-AGB star models that impressively reproduce a number of observational constraints (Pastorelli et al., 2019, 2020).
@@
Here, we take the next step in SFH modeling by combining the COLIBRI AGB star models with the well-tested CMD modeling framework MATCH (Dolphin, 2002) in order to measure quantitative SFHs exclusively from NIR observations of AGB stars. MATCH is a widely used, mature CMD fitting code that has been used to measure SFHs for galaxies from resolved star CMDs (e.g., Skillman et al., 2003; Williams et al., 2009; McQuinn et al., 2010; Monelli et al., 2010; Weisz et al., 2011; Cole et al., 2014; Weisz et al., 2014; Lewis et al., 2015; McQuinn et al., 2015; Skillman et al., 2017; Williams et al., 2017; Savino et al., 2023; Weisz et al., 2023a; McQuinn et al., 2024a, b, c). Though the incorporation of the COLIBRI models into MATCH is challenging (e.g., because of the need to correctly accounting for self-extinction in AGB stars), it provides several important benefits including a well-vetted framework for CMD analysis, robust uncertainty determination, and uniformity with MATCH-based SFHs of nearby galaxies based on optical CMDs, which we use to validate the reliability of our AGB star SFHs.
@@
Using this newly established framework, we take several steps to illustrate the technique in practice. We first fit mock data to illustrate self-consistency in our modeling. We then measure the SFH of M31’s outer disk using ground-based NIR observations. This region was selected to overlap with the Panchromatic Hubble Andromeda Treasury (PHAT) survey (Dalcanton et al., 2012b; Williams et al., 2014, 2023), which measured the spatially resolved SFH of the same region using deep Hubble Space Telescope (HST) optical CMDs (Williams et al., 2017, hereafter W17), providing a useful comparison point for our AGB star SFHs. Finally, we demonstrate the scientific potential of this new method by measuring the spatially resolved SFH of M31’s inner stellar halo using only ground-based NIR CMDs of AGB stars.
@@
Throughout this paper, we adopt the following M31 geometric parameters used by the PHAT team: center R.A. and Declination = (, ) and position angle ( (McConnachie et al., 2005; Barmby et al., 2006). We assume the inclination angle of M31 to be (Nieten et al., 2006). Unless otherwise noted, we use projected distances in M31.
@@ 2 AGB Theoretical Isochrones
The top panel shows AGB models with the number of AGB star thermal pulse cycles and the bottom panel shows AGB models with .
@@
Young AGB stars are typically the more luminous cool stars on the HR diagram, and have higher initial masses than older and fainter AGB stars. However, an important observational property is that even faint, lower-mass AGB stars are generally more luminous than the tip of the red giant branch (TRGB) in the near-IR. Moreover, the change in AGB age as a function of luminosity (i.e., the age gradient on the HR diagram) is much more pronounced for AGB stars than RGB stars.
@@
The age range we plot is for illustrative purposes only. AGB stars can exist down to Myr in age; below this age, stars do not progress through the AGB star phase because they are too massive to undergo the AGB phase. For the very oldest ages ( Gyr ago), age information comes primarily from E-AGB stars, as lower mass stars generally do not go through the TP-AGB phase. As shown in Figure 1, there is decreased separation between the RGB, E-AGB, and TP-AGB (when present) phases at older ages. In principle, it it possible to measure SFHs using only AGB stars back to Gyr ago, as demonstrated by our simulations in Appendix B.
Figure 2 shows a NIR CMD, containing stars with a variety of ages at a fixed metallicity of [M/H] dex. A modest amount of simulated noise has been added to the CMD to illustrate the AGB stars of different ages are readily distinguishable in a NIR CMD. Specifically, we generated and applied ASTs with a 50% completeness level at mag. The location of AGB stars in the color combination is highlighted by the grey box, which approximately denotes the location of the AGB stars as defined in the COLIBRI stellar isochrones. These same AGB stars are not easily distinguished from RGB stars at optical wavelengths, as we illustrate in later in the context of M31. Even in the red-optical (e.g., Å), the spectral energy distributions (SEDs) of AGB stars are largely indistinguishable from RGB stars; shifting to even bluer wavelengths makes separation of RGB and AGB stars less tenable. The m wavelength range appears optimal for isolating AGB stars on a CMD. Finally, we note that ‘extreme’ AGB stars (i.e., the reddest AGB stars with mag), which suffer from particularly complex physics (e.g., rapid and irregular mass-loss, circumstellar extinction), are so rare that they do not meaningfully contribute to the SFH determinations we describe in this paper.
In Figure 3 we show the behavior of AGB stars with a range of metallicities at a fixed age of 1 Gyr. Though there is clearly some dependence on metallicity, the influence of age is much more pronounced. For a deeper demonstration of this point, Figure 11 in Parada et al. (2021) shows COLIBRI isochrones with ages log(age) = 8.6 – 9.4 dex (i.e, 400 Myr – 2.5 Gyr) and metallicities (i.e., [M/H] dex). Above our cutoff selection magnitude for AGB stars of mag, the average magnitude of the youngest 400 Myr AGB stars is about mag, whereas the average magnitude of the oldest 2.5 Gyr AGB stars is about 1 mag fainter at mag. On the other hand, the most metal-poor AGB stars at [M/H] = -1.2 dex have an average magnitude of about mag, whereas the average magnitude of the metal-rich AGB stars at [M/H] = 0.0 dex is much more similar at mag. In summary: for AGB stars, a spread in log(age) of 0.8 dex spans mag, whereas a spread in metallicity of 1.2 dex spans mag, demonstrating the age of AGB stars affects their magnitudes more significantly than their metallicities.
@@ 3 Methodology
To fit the observed CMDs, we employ the widely-used CMD fitting package MATCH (version 2.7, Dolphin 2002, 2012, 2013). Given a set of stellar models, a specified initial mass function (IMF), and a unresolved binary fraction, MATCH creates and linearly combines simple stellar population (SSP) models to create a composite CMD. It then convolves the composite CMD with a noise model (i.e., bias, scatter, completeness) from ASTs to create a model CMD with the same observational effects as the data. The model and observed CMD are converted into Hess diagrams (i.e., a binned CMD) and compared using a Poisson likelihood function. MATCH then iterates through various weights of the SSP basis functions until the overall likelihood function is maximized. The synthetic model CMD that best matches the observed CMD is used as the best fit SFH. Random uncertainties on the SFH are computed using a Markov chain Monte Carlo approach, as we describe below.
@@
Starting from the COLIBRI models, we generate a grid of single stellar population models, each with a total stellar mass of (which yields AGB stars), spanning a range of ages and metallicities. Our SSP models have a metallicity range of [M/H] to 0.0 with a spacing of 0.2 dex, and ages from dex (100 Myr to 7.9 Gyr in linear time), logarithmically spaced by 0.1 dex. In principle, we could generate AGB stars models back to 14 Gyr ago for SFH determination. However, we stop at a younger lookback time in this study (1) because we were limited by the completeness of our datasets, and (2) to match the PHAT SFH we use to empirically validate our method; both reasons are described further in §5.2. We also note that the single stellar populations we generated are close enough to each other in log(age) and metallicity space that the artificial star based photometric errors would fill in the gaps between them. For example, a simple stellar population of log(age)=9.2 dex sufficiently represents the stellar population range from log(age)=9.15 to 9.25 dex.
@@
We generate the SSPs using a Kroupa IMF (Kroupa, 2001) and assume only single stars. The latter is generally a poor assumption of most parts of the CMD. However, the high luminosities of AGB stars, particularly in the NIR, mean that most binary combinations (e.g., AGB-MS star binary) do not significantly affect the colors and luminosities of AGB stars. However, an important caveat to this assumption is that our model ignores binaries with a surviving secondary star now on the AGB (e.g., an interacting system born with a primary and a secondary now on the AGB whose current mass is larger than its initial mass). The fraction of AGB stars that originate in binaries is unclear, and determining that fraction is currently beyond the scope of this paper.
@@
| Stellar Density | J | K | Number of | Region |
@@
| 25 | 18.7 | 17.7 | 504696 | Disk |
| Stellar Density | J | K | Number of | RegionsbbRegion borders shown in Figure 9. |
| 0 – 6 | 19.4 | 18.4 | 89530 | 1, 5 |
| 6 – 10 | 19.4 | 18.4 | 166270 | 4, 6 |
| 10 – 14 | 19.2 | 18.4 | 204640 | 2, 3 |